The following equations describe an economy. (Think of C, I, G, etc as being measured in billions and I as a percentage; a 5 percent interest rate implies i = 5.)

C = 0.8(1 – t)Y

t = 0.25

I = 900 – 50i

G = 800

Md/P = 0.25Y – 62.5i

Ms/P = 500


Derive the equations that describes the IS and LM curves?     

What is the equilibrium level of income and interest rate?  


1
Expert's answer
2021-11-23T11:05:57-0500

"a.\\
Y=C+I+G=0.8(1-t)Y+900-50R+800=0.6Y+900-50R+800=0.8Y+1700.8-50R\\

Y-0.6Y=1700-50R\\

0.4Y=1700-50R\\

Y=4250-125R"

"0.25Y-62.5R=500

0.25Y-500=62.5R\\

R=0.004Y-8\\

\\

LM=0.004Y-8"

"b.
\\Y=4250-125(0.004Y-8)\\

Y=4250-0.5Y+1000\\

Y+0.5Y=5250\\

1.5Y=5250\\

Y=3500\\



R=0.004\times 3500-8=6"